9x - [3x - (5x + 4)]
Answer by josgarithmetic(39613)

Parentheses first! Work from most inward and go outward from there.

9x-[3x-(5x+4)]=9x-[3x--5x-4], you can do nothing between the 5x and 4 because they are unlike terms. Notice the use of sign done in the subtraction of the expression.
9x-[3x--5x-4]=9x-[8x-4]
9x-[8x-4]=9x-8x+4, you subtract the entire bracketed expression, in this case accounting for sign and subtraction.
9x-8x+4=x+4, Additions from left to right, or if none, then subtractions from left to right.

Actually, you could change ALL subtractions into addition of the additive inverse but usually just dealing with subtraction by understanding and intuition is good enough.